OpenStudy (anonymous):

A rectangular storage container with an open top is to have a volume of 10 m3. The length of this base is twice the width. Material for the base costs $15 per square meter. Material for the sides costs $9 per square meter. Find the cost of materials for the cheapest such container. (Round your answer to the nearest cent.)

OpenStudy (anonymous):

All I have so far is: V = lwh ⇒ 10 = (2w)(w)h = 2w^2* h, so h = 5/w^2

OpenStudy (tamtoan):

now write the cost for each face @help_needed

OpenStudy (anonymous):

Its 2lw + 2wh + 2lh. So u mean 2(15) + 2(9) + 2(9) = 66.

OpenStudy (anonymous):

I thought u had to differentiate some function and idk what else. Its a calc problem

OpenStudy (tamtoan):

you were given only unit cost :) check again

OpenStudy (anonymous):

No its a calculus problem and i saw teh asnwer key and they differentiate a function and stuff. I just dont understand their explanation

OpenStudy (anonymous):

And 66 isnt the answer also. Its $163

OpenStudy (tamtoan):

yeah, that's why you have to multiply the unit cost by total area then you will get equation with variables so you can take derivative

OpenStudy (anonymous):

idk what u mean. Im honesltly bad at these calc problems. Will u be able to show work?

OpenStudy (tamtoan):

can you try to draw a picture for me ? i am bad at drawing :)

OpenStudy (anonymous):

ok. give me a sec

OpenStudy (tamtoan):

also lable each face of that box ...:)

OpenStudy (anonymous):

i hope thats good enuf for u

OpenStudy (anonymous):